3. Let a > 0) and consider the region E such that x > -a, y > 0, z > 0 and bounded by x + ay+a+z = a? (a) For a = 2 write outCould you also please explain why 'a' is significant here? Is it just suppose to symbolize as some constant? Can 'a' be anything?

Thank you much in advance, I will upvote.
Answer #1

27-a, 970,270 at ay tarz=a3 for a=2 Plane becomes x+2y + 42=8 - 7 2 + 2 2 = 1 The 110,0.1) To.4.0) SSS $dva 2122 12.0.6) I dqhere a is just a symbol. It can be any constant.